OpenStudy (anonymous):

tell whether the Mean Value theorem can be applied to f on the closed interval [a,b]. find all values of c in the open interval (a,b) such that f'(c)=f(b)-f(a)/b-a f(x)= (2-x)^(1/2), [-7,2]

OpenStudy (anonymous):

is \[f(x)=\sqrt{2-x}\] continuous on \([-7,2]\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hmm maybe not

OpenStudy (anonymous):

i mean the conclusion of the mvt might be true in any case, but i don't think this function is continuous on the closed interval, as it does not even exist to the right

OpenStudy (anonymous):

so then how can i show that on paper

OpenStudy (anonymous):

you can just say it

OpenStudy (anonymous):

the domain of this function is \((-\infty,2]\) but it is continuous from the left so that might be good enough. lets see if we can find that famous \(c\)

OpenStudy (anonymous):

we need some numbers \[f(-7)=\sqrt{2+7}=3\]\[f(-2)=0\] \[\frac{f(2)-f(-7)}{2-(-7)}=\frac{3}{9}=\frac{1}{3}\]

OpenStudy (anonymous):

messed that up didn't i?

OpenStudy (anonymous):

\[\frac{f(2)-f(-7)}{2-(-7)}=\frac{-3}{9}=-\frac{1}{3}\]

OpenStudy (anonymous):

i think so,

OpenStudy (anonymous):

and \[f'(x)=-\frac{1}{2\sqrt{2-x}}\]

OpenStudy (anonymous):

last job is to solve \[-\frac{1}{2\sqrt{2-x}}=-\frac{1}{3}\]

OpenStudy (anonymous):

i have a question in why would you put 3 on top, 2\[2-(-7)=9\]

OpenStudy (anonymous):

\[f(b)=f(2)=0\] \[f(a)=f(-7)=3\] \[f(b)-f(a)=f(2)-f(-7)=0-3=-3\]

OpenStudy (anonymous):

ok, so that is all you need to do to solve this

OpenStudy (anonymous):

it is numbers don't forget

OpenStudy (anonymous):

\[\frac{f(b)-f(a)}{b-a}\] is a number it is the slope of the secant line in this case that number is \(-\frac{1}{3}\)

OpenStudy (anonymous):

so then, that is what the question id asking for

OpenStudy (anonymous):

the second part of the question is asking "solve for \(c\) \[-\frac{1}{2\sqrt{2-c}}=-\frac{1}{3}\]

OpenStudy (anonymous):

then what would you plug in for c

OpenStudy (anonymous):

you don't plug in anything you solve that equation

OpenStudy (anonymous):

oh so then set it to zero to find c

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i think you are thinking way too hard solve for \(c\) you can probably do it in your head

OpenStudy (anonymous):

you got this?

OpenStudy (anonymous):

im lost, solve for c, would you do the derivative or cross multiply?

OpenStudy (anonymous):

i would start with \[2\sqrt{2-c}=3\]then square both sides

OpenStudy (anonymous):

get \[2-c=\frac{9}{4}\] then \[c=-\frac{1}{4}\]

OpenStudy (anonymous):

did you divide by two

OpenStudy (anonymous):

divide by two then square or square and then divide by 4 either way

OpenStudy (anonymous):

i was just about to say, did you square the square root

OpenStudy (anonymous):

yes square both sides

OpenStudy (anonymous):

lets write it all out with all the steps there is a \(c\in (-7,2)\) with \[f'(c)=\frac{f(2)-f(-7)}{2-(-7)}\] \[-\frac{1}{2\sqrt{2-c}}=-\frac{1}{3}\] \[2\sqrt{2-c}=3\] \[\sqrt{2-c}=\frac{3}{2}\] \[2-c=\frac{9}{4}\] \[c=2-\frac{9}{4}=-\frac{1}{4}\]

OpenStudy (anonymous):

ok thank you, im sorry it just didnt see it for some reason